OpenStudy (anonymous):

How can you tell whether vectors are linear independent? such as v_1 = (0, 0, 2, 2) v_2=(3, 3, 0, 0) v_3=(1,1,0,-1) v_4=(4, 4, 2, 1)

OpenStudy (anonymous):

A linearly DEPENDENT vector is a scalar multiple of another. So for example, consider: \[v=<1,1,1>\] If you had the vector <3,3,3> it would be: \[<3,3,3>=<3(1),3(1),3(1)>=3v\] So the vectors are linearly dependent. However, the vector <3,3,2> is linearly independent.

OpenStudy (anonymous):

A set of vectors \[B=\{v_1, v_2, v_3, ..., v_n\}\] if for \[a_1v_1 + a_2v_2 + ... + a_nv_n = 0\] only when \[a_i = 0\] for all 1 < i < n In other words no vector in the set can be expressed as a linear combination of the rest of the vectors.

OpenStudy (anonymous):

so the only way it can be linearly independent if you can't multiply any scalar to get the same vector

OpenStudy (anonymous):

so in my case it would be linearly independent?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

How come? You can't multiply all the vectors by a scalar to get all the same vector

OpenStudy (zarkon):

(0, 0, 2, 2) +(3, 3, 0, 0) +(1,1,0,-1) -(4, 4, 2, 1)=(0,0,0,0)

OpenStudy (zarkon):

Looks dependent to me

OpenStudy (anonymous):

slightly confused now so you're saying if we add them and get the zero vector it's dependent?

OpenStudy (anonymous):

How about in the case of V=<1,1> and v_1=<2,2>? Aren't they dependent but when added it's not a zero vector

OpenStudy (zarkon):

a nontrivial linear combination yields the zero vector...thus they are dependent

OpenStudy (zarkon):

2*(1,1)+(-1)(2,2)=(0,0)

OpenStudy (anonymous):

okay I think I get it now thanks!

OpenStudy (anonymous):

Mine is somewhat of a specific case. Using the combination of <1,1,1>, <1,1,1>, and <1,1,1> you can get <3,3,3> that I have above. Or treat it as a scalar multiple.